OpenStudy (anonymous):

Which is the best estimate for the quotient of 8,004.732 ÷ 0.196? A. 40 B. 400 C. 4,000 D. 40,000

OpenStudy (mathstudent55):

Round off each number. What does 8004.732 round off to?

OpenStudy (anonymous):

8000.000

OpenStudy (mathstudent55):

8000. good Now round off 0.196 to the nearest tenth.

OpenStudy (anonymous):

0.200

OpenStudy (mathstudent55):

Good. You are dividing 8000 by 0.2 Remember that a fraction also means division. \(2000 \div 0.2 = \dfrac{8000}{0.2} \) Ok?

OpenStudy (anonymous):

40000 and drop a 0 would make it 4000 C

OpenStudy (mathstudent55):

You can find equivalent fractions by multiplying the numerator and denominator by the same number. Let's multiply our denominator and numerator by 10 to eliminate the decimal. \(2000 \div 0.2 = \dfrac{8000}{0.2} = \dfrac{8000 \times 10}{0.2 \times 10} = \dfrac{80,000}{2}\) Now youj can do the division easily. What is \(80,000 \div 2\) ?

OpenStudy (mathstudent55):

Sorry, above I meant to write \(8000 \div 0.2 = \dfrac{8000}{0.2} \)

OpenStudy (mathstudent55):

The answer is 40,000. Don't drop a zero.